A steel ball weighing 128 Ib is suspended from a spring, whereupon the spring is stretched 2 ft from its natural length. The ball is started in motion with no initial velocity by displacing it 6 in above the equilibrium position. Assuming no air resistance, find (a) an expression for the position of the ball at any time t, and (b) the position of the ball at t=pi/4 sec
1
Expert's answer
2021-10-11T15:53:19-0400

Steel ball weight=128lb

∆x=2ft

At equilibrium x=0

Position above equilibrium, t=0,x=6in=0.5ft

x=ACos wt="\frac{1}{2}Cos (\frac{2πt}{T})"

f=ma=-kx

"\frac{dx}{dt}=\frac{d}{dt}(\frac{1}{2}Cos(\frac{2πt}{T}))"

"=\frac{2π}{2T}(-Sin\frac{2πt}{T})"

a="\frac{d^{2}x}{dt^{2}}=\frac{d}{dt}(-\frac{2π}{2T}Sin\frac{2πt}{T})"

"=\frac{1}{2}(\frac{2π}{T})^{2}Cos \frac{2πt}{T}"

"K=\frac{F}{∆x}" ="\frac{128}{2}=64lb/ft"

m="\frac{128}{32}=4"

a)

"x=\frac{1}{2}Cos (\frac{2πt}{\frac{π}{2}})"

"=0.5 Cos(4t)"

b)

t="\frac{π}{4}"

"x=0.5Cos(4*\frac{π}{4})"

=0.5Cosπ

=-0.5